Circular No. 7820

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ap IN M74 S. Mattila and P. Meikle, Imperial College, London, report that SN

2002ap continues to brighten in the infrared, as indicated by the following magnitudes from observations obtained by K. Chambers (Institute for Astronomy, University of Hawaii) at the U.K. Infrared Telescope (+ infrared imager UFTI): Feb. 3.2 UT, I = 12.62, Z = 12.44, J = 12.39, H = 12.23, K = 12.20; 4.2, 12.51, 12.30, 12.27, 12.08, 11.94.

L. Wang, Lawrence Berkeley National Laboratory; D. Baade, European Southern Observatory (ESO); C. Fransson, Stockholm Observatory; P. Hoeflich, University of Texas; P. Lundqvist, Stockholm Observatory; and J. C. Wheeler, University of Texas, report that high-quality spectropolarimetry (range 350–860 nm; spectral resolution 1.27 nm and 0.265 nm/pixel) of SN 2002ap was obtained with the ESO Very Large Telescope UT3 (+ FORS1) on Feb. 3 UT. A broad spectral absorption feature at 650– 800 nm (which, if due to the Ca II infrared triplet, implies an expansion velocity of order 45 000 km/s) shows a degree of polarization distinctively different from other parts of the observed spectrum. The degree of polarization of this feature peaks at 735.1 nm (with FWHM = 64.3 nm) and differs by > 1.1 percent when compared to neighboring wavelength regions. As the contribution from interstellar dust is uncertain, the observed 1.1percent difference should be regarded as a lower limit to the polarization intrinsic to the supernova and implies the presence of significantly asymmetric ejecta. Further spectropolarimetric observations are encouraged.

P. Vreeswijk, University of Amsterdam; and S. Smartt, Institute of Astronomy, Cambridge, report that images taken by G. Gilmore and R. Wyse on the Isaac Newton Telescope at La Palma on Feb. 2.9 UT have allowed an accurate projection of the position of SN 2002ap onto the prediscovery images reported on IAUC 7816: “Within the new errors, the supernova is not coincident with the V = 21.2 object reported as a possible progenitor. The difference in position between SN 2002ap and this object is 1′′.7 ± 0′′.3, where the error is dominated by the error in the position of the faint object, indicating that this was a chance alignment. There is no object in the prediscovery images at the supernova position, to limiting (5σ, 2′′ aperture radius) magnitudes of U = 20.8, B = 22.1, V = 22.0, I = 20.8.”

W. Hasubick, Buchloe, Germany, reports that CCD frames taken on Jan. 4.795 UT show nothing at the position of SN 2002ap (limiting red mag 18.5). CCD R magnitudes by K. Hornoch, Lelekovice, Czech Republic: Jan. 31.808 UT, 13.04; Feb. 2.764, 12.66; 3.749, 12.53.

2002 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green